# Branching Brownian motion versus Random Energy Model in the supercritical phase: overlap distribution and temperature susceptibility

### v2

@@ 4.3 Comparison wih the REM in the general decorated case
Note that, for this first family of examples, the difference $\kappa_{d}(\beta)-\kappa(\beta)$ grows at most like $(\beta-1)^{-2}$, which is the order of magnitude of $\kappa(\beta)$ as $\beta\downarrow 1$. More precisely, if $b\leq 3$, then $\kappa_{d}(\beta)\sim\kappa(\beta)$, and if $b>3$, then both susceptibilities grow at the same speed but with different multiplicative constants.
In the case $b>3$, observe that the expectations in Equation (4.10) are all dominated by the event where $X\asymp(\beta-1)^{-1}$, so, on this event, $S_{\beta}^{\prime}/S_{\beta}\asymp(\beta-1)^{-1}$ which yields eventually a variance $\operatorname{Var}_{\beta}(S_{\beta}^{\prime}/S_{\beta})$ of order $(\beta-1)^{-2}$. This observation is the idea motivating the next family of examples: in order to construct cases where $\kappa_{d}(\beta)$ grows much faster than $\kappa(\beta)$, we want the expectations in Equation (4.10) to be dominated by the event where $X\asymp(\beta-1)^{-\alpha}$ for some $\alpha>1$.
@@
$\kappa_{d}(\beta)-\kappa(\beta)\sim\frac{3\gamma^{\frac{1}{1-\gamma}}}{2\pi^{2}(1-\gamma)}(\beta-1)^{-\frac{2-\gamma}{1-\gamma}}.$
Note that the exponent $\frac{2-\gamma}{1-\gamma}$ is always larger than 2, so that $\kappa_{d}(\beta)$ grows faster than $\kappa(\beta)$, and this exponent can be made arbitrarily large by taking $\gamma$ close to 1.
@@ 4.4 The BBM case
We first prove that $\kappa_{d}(\beta)>\kappa(\beta)$, when $\beta>1$ is fixed. By Proposition 4.5, it is sufficient to show that
@@
This event has positive $\mathbb{P}$-probability by [12, Proposition 3.4]. So, on this event intersected with the one where Equation (4.16) holds (this intersection being non-empty), we have
@@
This is a slightly stronger version of [21, Lemma 5.2] where no uniformity for $v\in[\theta\sqrt{r},\theta^{-1}\sqrt{r}]$ is stated: however, the aforementioned claim follows from their proof (more precisely, it follows from [21, Lemma 5.6] in the same way as Lemma 4.10 below follows from Lemma A.1). Then, we deduce from this claim that, for $\beta>1$ close enough to 1, up to a modification of $t_{0}$, we also have, for any $t\geq t_{0}$,
@@
We proceed by contradiction: assume that, for any $a>0$ and $0<b<B$, $\kappa=\kappa^{\prime}$. Then, differentiating the relation $\kappa=\kappa^{\prime}$ w.r.t. $B$, we get that, for any $a,B>0$, the following quantity is the same for $k=1$ and $k=2$:
@@ A.1 Uniform bounds for moments of level sets
The bound on the second moment is a direct consequence of Lemma A.6, so we focus here on the first moment. However, it could also be deduced from the proof of [21, Lemma 5.3], in a similar way as what is done below for the first moment.
To bound uniformly $\widetilde{\mathbb{E}}_{t}[\mathcal{C}_{t,r_{t}}^{*}([-v,0])]$, we follow the proof of [21, Lemma 5.2], which establishes an asymptotic equivalent for the first moment of $\mathcal{C}_{t,r_{t}}^{*}([-v,0])$ as $t\to\infty$ and then $v\to\infty$. The proof begins by writing
@@
by [21, Lemmata 3.1 and 3.4] (the constant $C_{1}$ equals $2f^{(0)}(0)g(0)$ with notation of [21, Lemma 3.4]). In particular, there exists a constant $c>0$ such that the probability in Equation (A.2) is at least $c/t$ for any $t\geq 1$. Therefore, it remains to prove that there exist $C>0$ and $t_{0}\geq 0$ such that, for any $t\geq t_{0}$ and $v\geq 0$,
@@
We now focus on proving Equation (A.3). Applying successively [21, Lemmata 5.4 and 5.5] (note that $j_{t,v}(s)=j_{t,v}^{\geq 0}(s)$), we get for any $t$ such that $r_{t}\leq t/2$ and any $v\geq 0$,
@@
This last integral is smaller than $C((v+1)^{-1}+\mathrm{e}^{-v/2})$ for any $v\geq 0$ (see [21, Equations (5.44)-(5.45)] for details) and so we get Equation (A.3). ∎
@@ A.2 First moment of level sets on a particular event
We prove in this subsection Lemma 4.10. This is a (non-trivial) refinement of the proof of [21, Lemma 5.2].
@@
which is analog to $B_{r,t}$ (see Equation (4.27)) but for the process $(\widehat{W}_{t,s})_{s\in[0,t]}$ defined in [21, Equation (3.1)] as
@@
where $J_{t,v}(s)$ is introduced in [21, Equation (5.9)], it follows from the proof of [21, Lemma 5.4], that
@@
As in [21, Equation (5.17)], for any $M\geq 0$, we split $j_{t,v,r}(s)$ into
@@
where the two last terms are negligible. Indeed, it is proved in [21, Equation (5.41)] (and the paragraph around) that
@@
for any $M>0$, $40\leq r\leq t/$ 4, $s\in[(1-\delta)r,(1+\delta)r]$ and $v\in[\theta\sqrt{r},\theta^{-1}\sqrt{r}]$ by [21, Lemma 5.5]. Finally, using Lemma A.1, we have, as $t\to\infty$, then $r\to\infty$ and then $M\to\infty$, uniformly in $v\in[\theta\sqrt{r},\theta^{-1}\sqrt{r}]$,
@@
Using the relation666 The authors of [21] do not keep precisely track of constants but this relation can be deduced from a careful reading of their paper. Alternatively letting $b\to 0$ and $B\to\infty$, we have $\mathbbm{1}_{B_{r,t}}\to 1$ and $\varphi_{b,B}\to 1$ and Equation (A.9) gives an alternative proof of their [21, Lemma 5.2], showing that the constant $C$ appearing there equals $2\sqrt{2\pi}C_{2}/C_{1}$. Then, a quick look at the proof of [21, Proposition 1.5] ensures that this constant $C$ is actually $C_{\star}$. $C_{\star}=2\sqrt{2\pi}\,C_{2}/C_{1}$ gives the desired result. ∎
Lemma A.1.
@@
where $f^{(0)},f,g$ are positive functions introduced in [21, Lemma 3.4].
@@
We follow ideas from the proof of [21, Lemma 5.6], but instead of only distinguishing according to the value of $\widehat{W}_{t,s}$, we also distinguish according to the value of $\widehat{W}_{t,r}$. For comparison, the constant $C$ appearing in the statement of [21, Lemma 5.6] equals $C_{2}$ introduced here.
@@
as $t\to\infty$ and then $r\to\infty$, uniformly in $s\in[(1+\delta)r,\delta^{-1}r]$, $y\in[-B\sqrt{r},-b\sqrt{r}]$ and $z\in[-M,M]$. Furthermore, it follows from [21, Lemma 3.4] that
@@
as $t\to\infty$ and then $r\to\infty$, uniformly in $s$, $y$ and $z$ as before. Finally, it follows from [21, Lemma 4.2] (see also [21, Equation (5.37)]) that
@@
as $t\to\infty$ and then $r\to\infty$, uniformly in $s$ and $v$ as before. A change of variable in the first integral and letting $M\to\infty$ in the second integral (the fact that this integral converges to a finite limit is justified at the end of the proof of [21, Lemma 5.6]) yields the result.
@@ A.3 Cross-moments of level sets
Our aim in this section is to prove Proposition 2.14. For this, we follow the proof of [21, Proposition 1.5], which bounds $\mathbb{E}[\mathcal{C}([-v,0])^{2}]$. This proof is based on a series of five lemmas, that we re-state here in a new version tuned for our purpose of dealing with two level sets of different levels $v$ and $v^{\prime}$.
In the following lemma, as in [21, Lemma 4.3], we work with a 2-spine BBM defined under the probability measure ~$\widetilde{\mathbb{P}}^{(2)}$ as follows. It starts with one particle at 0 at time 0 which is part of the spines 1 and 2. Particles belonging to $m$ spines branch at rate $2^{m}$ and move according to a standard Brownian motion. At a branching point, for each $k\in\{1,2\}$, if the parent was part of spine $k$, then one of both children is chosen uniformly at random to be part of spine $k$. We denote $X_{t}(k)$ the particle at time $t$ which is part of the spine $k$.y For $x,y\in L_{t}$, we write $d(x,y)=r$ if the most recent common ancestor of $x$ and $y$ died at time $t-r$.
Lemma A.2.
@@
This is a new version of [21, Lemma 4.3] and we explain how to adapt its proof. Similarly as [21, Equation (4.15)], the probability in the statement equals
@@
For $z>u$, using [21, Equation (4.2)] for the first one and [21, Equation (4.3)] for the second one, we bound the product of the two first probabilities in Equation (A.10) by
@@
This is exactly the same as [21, Equation (4.20)] up to the factor $(u-v^{\prime}+1)\mathrm{e}^{-\sqrt{2}v^{\prime}}$ where primes have been added. Therefore, this part of the integral is dealt with exactly the same way as in [21].
For $z\leq u$, we use [21, Equation (4.2)] for both first probabilities in Equation (A.10) and note that, because $v\leq v^{\prime}$,
@@
There are two differences with [21, Equation (4.19)]: the factor $(u-v^{\prime}+1)\mathrm{e}^{-\sqrt{2}v^{\prime}}$ where primes have been added (but this adds no new difficulty), and the additional term $\mathrm{e}^{(v-z)/2}$ in the last parentheses. This latter gives rise to the following new term, which should be added to the integral in [21, Equation (4.21)],
@@
which can be included in the upper bound of the statement of the lemma after taking care of the other factors (see [21, Equation (4.22)]: this is exactly how another part of the integral in [21, Equation (4.21)] is bounded). ∎
Lemma A.3.
@@
This follows from Lemma A.2 in the same way as [21, Lemma 4.4] follows from [21, Lemma 4.3]. ∎
Lemma A.4.
@@
This follows directly from the proof of [21, Lemma 5.4]. ∎
Lemma A.5.
@@
This is a new version of [21, Lemma 5.5].
@@
The case $s=s^{\prime}$ is also identical to the proof of [21, Lemma 5.5], applying here Lemma A.3 instead of [21, Lemma 4.4]. ∎
Lemma A.6.
@@
by integrating w.r.t. $s^{\prime}$ first. Then, proceeding as in the proof of [21, Lemma 5.3], this last integral is at most $C/(v+1)$ and this yields the result. ∎
@@
This follows from Lemma A.6 in the same way as [21, Proposition 1.5] follows from [21, Lemma 5.3]. ∎
@@ A.4 Small moments of $S_{\beta}$
We first decompose $\mathcal{C}_{t,r_{t}}^{*}(f_{\beta})$ along the spine as it is done for level sets in [21, Lemma 5.4]. Setting (this replaces $J_{t,v}(s)$ defined in [21, Equation (5.9)])
@@
where $\mathcal{N}$ is a Poisson point process on $\mathbb{R}_{+}$ with intensity $2\mathop{}\mathopen{}\mathrm{d}s$. Using subadditivity of $x\mapsto x^{\gamma}$ (note that the integral above is actually a finite sum) and then proceeding as in the proof of [21, Lemma 5.4], we get
@@
For this, we follow the ideas from the proof of [21, Lemma 5.5] (in the case $M=0$). Conditioning on $\widehat{W}_{t,s}$, we get, similarly as [21, Equation (5.20)],
@@
Using [21, Lemma 4.2], the right-hand side of Equation (A.16) is at most
@@
Coming back to Equation (A.15) and bounding the other factors in the integrand as in [21, Equations (5.23) and (5.24)], we get

### v3

@@ 4.3 Comparison wih the REM in the general decorated case
Note that, for this first family of examples, the difference $\kappa_{\mathrm{d}}(\beta)-\kappa(\beta)$ grows at most like $(\beta-1)^{-2}$, which is the order of magnitude of $\kappa(\beta)$ as $\beta\downarrow 1$. More precisely, if $b\leq 3$, then $\kappa_{\mathrm{d}}(\beta)\sim\kappa(\beta)$, and if $b>3$, then both susceptibilities grow at the same speed but with different multiplicative constants.
In the case $b>3$, observe that the expectations in Equation (4.10) are all dominated by the event where $X\asymp(\beta-1)^{-1}$, so, on this event, $S_{\beta}^{\prime}/S_{\beta}\asymp(\beta-1)^{-1}$ which yields eventually a variance $\operatorname{Var}_{\beta}(S_{\beta}^{\prime}/S_{\beta})$ of order $(\beta-1)^{-2}$. This observation is the idea motivating the next family of examples: in order to construct cases where $\kappa_{\mathrm{d}}(\beta)$ grows much faster than $\kappa(\beta)$, we want the expectations in Equation (4.10) to be dominated by the event where $X\asymp(\beta-1)^{-\alpha}$ for some $\alpha>1$.
@@
$\kappa_{\mathrm{d}}(\beta)-\kappa(\beta)\sim\frac{3\gamma^{\frac{1}{1-\gamma}}}{2\pi^{2}(1-\gamma)}(\beta-1)^{-\frac{2-\gamma}{1-\gamma}}.$
Note that the exponent $\frac{2-\gamma}{1-\gamma}$ is always larger than 2, so that $\kappa_{\mathrm{d}}(\beta)$ grows faster than $\kappa(\beta)$, and this exponent can be made arbitrarily large by taking $\gamma$ close to 1.
@@ 4.4 The BBM case
We first prove that $\kappa_{\mathrm{d}}(\beta)>\kappa(\beta)$, when $\beta>1$ is fixed. By Proposition 4.5, it is sufficient to show that
@@
This event has positive $\mathbb{P}$-probability by [14, Proposition 3.4]. So, on this event intersected with the one where Equation (4.16) holds (this intersection being non-empty), we have
@@
This is a slightly stronger version of [26, Lemma 5.2] where no uniformity for $v\in[\theta\sqrt{r},\theta^{-1}\sqrt{r}]$ is stated: however, the aforementioned claim follows from their proof (more precisely, it follows from [26, Lemma 5.6] in the same way as Lemma 4.10 below follows from Lemma A.1). Then, we deduce from this claim that, for $\beta>1$ close enough to 1, up to a modification of $t_{0}$, we also have, for any $t\geq t_{0}$,
@@
We proceed by contradiction: assume that, for any $a>0$ and $0<b<B$, $\kappa=\kappa^{\prime}$. Then, differentiating the relation $\kappa=\kappa^{\prime}$ w.r.t. $B$, we get, for any $a,B>0$, the following quantity is the same for $k=1$ and $k=2$:
@@ A.1 Uniform bounds for moments of level sets
The bound on the second moment is a direct consequence of Lemma A.6, so we focus here on the first moment. However, it could also be deduced from the proof of [26, Lemma 5.3], in a similar way as what is done below for the first moment.
To bound uniformly $\widetilde{\mathbb{E}}_{t}[\mathcal{C}_{t,r_{t}}^{*}([-v,0])]$, we follow the proof of [26, Lemma 5.2], which establishes an asymptotic equivalent for the first moment of $\mathcal{C}_{t,r_{t}}^{*}([-v,0])$ as $t\to\infty$ and then $v\to\infty$. The proof begins by writing
@@
by [26, Lemmata 3.1 and 3.4] (the constant $C_{1}$ equals $2f^{(0)}(0)g(0)$ with notation of [26, Lemma 3.4]). In particular, there exists a constant $c>0$ such that the probability in Equation (A.2) is at least $c/t$ for any $t\geq 1$. Therefore, it remains to prove that there exist $C>0$ and $t_{0}\geq 0$ such that, for any $t\geq t_{0}$ and $v\geq 0$,
@@
We now focus on proving Equation (A.3). Applying successively [26, Lemmata 5.4 and 5.5] (note that $j_{t,v}(s)=j_{t,v}^{\geq 0}(s)$), we get for any $t$ such that $r_{t}\leq t/2$ and any $v\geq 0$,
@@
This last integral is smaller than $C((v+1)^{-1}+\mathrm{e}^{-v/2})$ for any $v\geq 0$ (see [26, Equations (5.44)-(5.45)] for details) and so we get Equation (A.3). ∎
@@ A.2 First moment of level sets on a particular event
We prove in this subsection Lemma 4.10. This is a (non-trivial) refinement of the proof of [26, Lemma 5.2].
@@
which is analog to $B_{r,t}$ (see Equation (4.27)) but for the process $(\widehat{W}_{t,s})_{s\in[0,t]}$ defined in [26, Equation (3.1)] as
@@
where $J_{t,v}(s)$ is introduced in [26, Equation (5.9)], it follows from the proof of [26, Lemma 5.4], that
@@
As in [26, Equation (5.17)], for any $M\geq 0$, we split $j_{t,v,r}(s)$ into
@@
where the two last terms are negligible. Indeed, it is proved in [26, Equation (5.41)] (and the paragraph around) that
@@
for any $M>0$, $40\leq r\leq t/$ 4, $s\in[(1-\delta)r,(1+\delta)r]$ and $v\in[\theta\sqrt{r},\theta^{-1}\sqrt{r}]$ by [26, Lemma 5.5]. Finally, using Lemma A.1, we have, as $t\to\infty$, then $r\to\infty$ and then $M\to\infty$, uniformly in $v\in[\theta\sqrt{r},\theta^{-1}\sqrt{r}]$,
@@
Using the relation777 The authors of [26] do not keep precisely track of constants but this relation can be deduced from a careful reading of their paper. Alternatively letting $b\to 0$ and $B\to\infty$, we have $\mathbbm{1}_{B_{r,t}}\to 1$ and $\varphi_{b,B}\to 1$ and Equation (A.9) gives an alternative proof of their [26, Lemma 5.2], showing that the constant $C$ appearing there equals $2\sqrt{2\pi}C_{2}/C_{1}$. Then, a quick look at the proof of [26, Proposition 1.5] ensures that this constant $C$ is actually $C_{\star}$. $C_{\star}=2\sqrt{2\pi}\,C_{2}/C_{1}$ gives the desired result. ∎
Lemma A.1 (BBM case).
@@
where $f^{(0)},f,g$ are positive functions introduced in [26, Lemma 3.4].
@@
We follow ideas from the proof of [26, Lemma 5.6], but instead of only distinguishing according to the value of $\widehat{W}_{t,s}$, we also distinguish according to the value of $\widehat{W}_{t,r}$. For comparison, the constant $C$ appearing in the statement of [26, Lemma 5.6] equals $C_{2}$ introduced here.
@@
as $t\to\infty$ and then $r\to\infty$, uniformly in $s\in[(1+\delta)r,\delta^{-1}r]$, $y\in[-B\sqrt{r},-b\sqrt{r}]$ and $z\in[-M,M]$. Furthermore, it follows from [26, Lemma 3.4] that
@@
as $t\to\infty$ and then $r\to\infty$, uniformly in $s$, $y$ and $z$ as before. Finally, it follows from [26, Lemma 4.2] (see also [26, Equation (5.37)]) that
@@
as $t\to\infty$ and then $r\to\infty$, uniformly in $s$ and $v$ as before. A change of variable in the first integral and letting $M\to\infty$ in the second integral (the fact that this integral converges to a finite limit is justified at the end of the proof of [26, Lemma 5.6]) yields the result.
@@ A.3 Cross-moments of level sets
Our aim in this section is to prove Proposition 2.15. For this, we follow the proof of [26, Proposition 1.5], which bounds $\mathbb{E}[\mathcal{C}([-v,0])^{2}]$. This proof is based on a series of five lemmas, that we re-state here in a new version tuned for our purpose of dealing with two level sets of different levels $v$ and $v^{\prime}$.
In the following lemma, as in [26, Lemma 4.3], we work with a 2-spine BBM defined under the probability measure ~$\widetilde{\mathbb{P}}^{(2)}$ as follows. It starts with one particle at 0 at time 0 which is part of the spines 1 and 2. Particles belonging to $m$ spines branch at rate $2^{m}$ and move according to a standard Brownian motion. At a branching point, for each $k\in\{1,2\}$, if the parent was part of spine $k$, then one of both children is chosen uniformly at random to be part of spine $k$. We denote $X_{t}(k)$ the particle at time $t$ which is part of the spine $k$. For $x,y\in L_{t}$, we write $d(x,y)=r$ if the most recent common ancestor of $x$ and $y$ died at time $t-r$.
Lemma A.2 (BBM case).
@@
This is a new version of [26, Lemma 4.3] and we explain how to adapt its proof. Similarly as [26, Equation (4.15)], the probability in the statement equals
@@
For $z>u$, using [26, Equation (4.2)] for the first one and [26, Equation (4.3)] for the second one, we bound the product of the two first probabilities in Equation (A.10) by
@@
This is exactly the same as [26, Equation (4.20)] up to the factor $(u-v^{\prime}+1)\mathrm{e}^{-\sqrt{2}v^{\prime}}$ where primes have been added. Therefore, this part of the integral is dealt with exactly the same way as in [26].
For $z\leq u$, we use [26, Equation (4.2)] for both first probabilities in Equation (A.10) and note that, because $v\leq v^{\prime}$,
@@
There are two differences with [26, Equation (4.19)]: the factor $(u-v^{\prime}+1)\mathrm{e}^{-\sqrt{2}v^{\prime}}$ where primes have been added (but this adds no new difficulty), and the additional term $\mathrm{e}^{(v-z)/2}$ in the last parentheses. This latter gives rise to the following new term, which should be added to the integral in [26, Equation (4.21)],
@@
which can be included in the upper bound of the statement of the lemma after taking care of the other factors (see [26, Equation (4.22)]: this is exactly how another part of the integral in [26, Equation (4.21)] is bounded). ∎
Lemma A.3 (BBM case).
@@
This follows from Lemma A.2 in the same way as [26, Lemma 4.4] follows from [26, Lemma 4.3]. ∎
Lemma A.4 (BBM case).
@@
This follows directly from the proof of [26, Lemma 5.4]. ∎
Lemma A.5 (BBM case).
@@
This is a new version of [26, Lemma 5.5].
@@
The case $s=s^{\prime}$ is also identical to the proof of [26, Lemma 5.5], applying here Lemma A.3 instead of [26, Lemma 4.4]. ∎
Lemma A.6 (BBM case).
@@
by integrating w.r.t. $s^{\prime}$ first. Then, proceeding as in the proof of [26, Lemma 5.3], this last integral is at most $C/(v+1)$ and this yields the result. ∎
@@
This follows from Lemma A.6 in the same way as [26, Proposition 1.5] follows from [26, Lemma 5.3]. ∎
@@ A.4 Small moments of $S_{\beta}$
We first decompose $\mathcal{C}_{t,r_{t}}^{*}(f_{\beta})$ along the spine as it is done for level sets in [26, Lemma 5.4]. Setting (this replaces $J_{t,v}(s)$ defined in [26, Equation (5.9)])
@@
where $\mathcal{N}$ is a Poisson point process on $\mathbb{R}_{+}$ with intensity $2\mathop{}\mathopen{}\mathrm{d}s$. Using subadditivity of $x\mapsto x^{\gamma}$ (note that the integral above is actually a finite sum) and then proceeding as in the proof of [26, Lemma 5.4], we get
@@
For this, we follow the ideas from the proof of [26, Lemma 5.5] (in the case $M=0$). Conditioning on $\widehat{W}_{t,s}$, we get, similarly as [26, Equation (5.20)],
@@
Using [26, Lemma 4.2], the right-hand side of Equation (A.16) is at most
@@
Coming back to Equation (A.15) and bounding the other factors in the integrand as in [26, Equations (5.23) and (5.24)], we get